A 200 N wagon is to be pulled up a 30 degrees incline at a constant speed. How large a force in kN parallel to the incline is needed if friction effects are negligible?


1
Expert's answer
2021-10-20T10:42:34-0400

Let's apply the Newton's Second Law of Motion:


"F_{pull}-Wsin\theta=0,""F_{pull}=Wsin\theta=200\ N\times sin30^{\circ}=100\ N=0.1\ kN."